A store has 100 bikes, they have 150 bikes rental per hour . The average bike rental lasts for 15 min . Rental fee for one bike is 10 cents per minute

1. the average number of bikes being rented is ? 2. This store's daily revenue is ? /day
Answer #1

Flow rate = 150 bikes per hour
Flow time = 15 minutes = 15/60 hours

From Little's law: WIP = Flow time x Flow rate = (15/60) x 150 = 37.5 bikes

So, 37.5 bikes are being rented at a time on an average.

Daily revenue

= 150 bikes per hour x 8 hours per day x 15 minutes per bike x 10 cents per minute

= 180,000 cents per day

= $1,800 per day (assuming that one day consists of 8 operating hours)